Asha and Samuel have bookshelves of the same size partly filled with books. Asha’s shelf is 5 / 6 th full and Samuel’s shelf is 2/ 5 th full. Whose bookshelf is more full? By what fraction? Solutions: Fraction of Asha’s bookshelf = 5 / 6 Fraction of Samuel’s bookshelf = 2 / 5 Convert these fractions into like fractions

Grade:12

1 Answers

Harshit Singh
3 years ago
Dear Student

5 / 6 = 5 / 6×5 / 5
= (5 × 5) / (6 × 5)
= 25 / 30
2 / 5 = 2 / 5 ×6 / 6
= (2 × 6) / (5 × 6)
= 12 / 30
25 / 30 > 12 / 30
5 /6 > 2 /5
∴Asha’s bookshelf is more full than Samuel’sbookshelf
Difference = 5 / 6–2 / 5
= 25 / 30–12 / 30
= 13 / 30

Thanks